# RadarPLM: Adapting Pretrained Language Models for Marine Radar Target Detection with Preference-aware Loss

### v3

@@ III-E Autoencoder-based Binary Classification Head Retraining
As shown in Fig. 1 and Section III-A, the PLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation. To achieve consistent decision-making, we retrain a binary classifier head using the learned representation $F^{\mathrm{LN}}$. An autoencoder-based biary classification head extracts informative features from this high-dimensional input for robust detection. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@
3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of reference model, see Eq.(12).
@@
    Fine-tune PLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(15).
@@
9:Fine-tuned RadarLLM model with controllable FAR.
@@ IV Experiments
Epochs for LLM fine-tuning 500
@@
TABLE III: Hyperparameters for Autoencoder network.
Block Layer Filter Stride Padding Output
[…]
Conv6+ReLU 16,1*1 1 Same 55*16
Flatten+Fully-Connected1 - - - 64
Fully-Connected2 - - - 20
Decoder Fully-Connected3+Reshape - - - 55*16
Conv1+ReLU 32,1*1 1 1 55*32
Conv2+ReLU 64,1*1 1 1 55*64
Conv3+ReLU 128,1*1 1 1 55*128
Conv4+ReLU 256,1*1 1 1 55*256
Conv5+ReLU 512,1*1 1 1 55*512
Conv6+ReLU 768,1*1 1 1 55*768
Classifier GeLU+LayerNorm - - - 20
Fully-Connected4 - - - 2
@@ IV-A4 Implementation Details
The experimentally relevant hyperparameters and details are shown in Table II. For patching, the size of feature patches is set to 48. For the pre-trained LLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed. The network hyperparameters of the Autoencoder are shown in Table III. The value of $\alpha$ in Eq.(14) is set to 0.9. All signal feature extraction operations are performed with MATLAB R2016a software. All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
IV-B Performance Comparison on Sufficient Training Samples
@@
Figure 6: Comparison of model detection performance on various datasets.
TABLE IV: Comparison of detection rates (%) on sufficient training samples.
Method $H$ HV VH $V$ AVG
RNN [14] 46.04 46.66 38.18 17.04 36.98
Bi-LSTM [14] 70.39 73.15 76.48 60.89 72.48
GRU [41] 76.46 77.55 73.96 67.48 73.86
Transformer [36] 52.95 71.73 67.32 55.67 61.92
PatchTST [42] 75.23 75.72 77.49 66.59 73.76
ResNet18 [10] 66.24 80.96 82.31 73.97 75.87
OFA [18] 74.19 80.38 79.71 70.36 76.16
ADN18 [43] 79.74 83.50 82.01 78.02 80.82
RadarPLM 82.51 84.81 84.61 75.90 81.96
(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
TABLE V: Network parameters (training parameters / total parameters) and average inference time of different methods.
@@
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure LABEL:DT presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
@@
Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarPLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. In particular, RadarPLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
IV-C Performance Comparison on Limited Training Samples
@@
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarPLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
In Table VI, we report the detection performance of RadarPLM and several baseline methods under scenarios with limited training data. Compared with the experimental setup described in Section IV-A1, the advantages of RadarPLM become more pronounced. Specifically, RadarPLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14]. This marked improvement can be mainly attributed to three key factors:
@@
TABLE VI: Comparison of detection rates (%) on limited training samples.
Method $H$ HV VH $V$ AVG
RNN [14] 35.49 30.13 31.82 15.14 28.15
Bi-LSTM [14] 61.96 67.46 65.39 42.19 59.25
GRU [41] 70.43 74.14 74.20 59.47 69.56
Transformer [36] 70.22 64.56 65.31 51.15 62.81
PatchTST [42] 68.31 72.30 75.69 61.77 69.52
ResNet18 [10] 63.71 78.33 78.08 71.05 72.79
OFA [18] 73.64 76.81 77.44 63.33 72.81
ADN18 [43] 75.45 81.50 80.81 70.97 77.18
RadarPLM 79.26 82.31 81.51 72.54 78.91
(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
(1). Data1
(2). Data2
(3). Data3
Figure 7: Comparison of detection rates between the proposed preference-aware loss and the conventional cross-entropy loss. We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
IV-D Ablation Study
IV-D1 Backbone Architectures
  1. 1.
  2. 2.
  3. 3.
In Table VII, we show the results of the ablation study on backbone choice. It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
TABLE VII: Results of ablation study on the backbone choice and fine-tuning strategies. RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results.
@@ IV-D1 Backbone Architectures
IV-D2 Fine-tuning Strategies
To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table VIII, and the corresponding results are presented in Table VII. In this experiment, four fine-tuning configurations were evaluated: (1) RadarPLM: jointly fine-tuning the LoRA modules and Layer Normalization (LN) parameters, (2) RadarPLM (LoRA): fine-tuning only the LoRA modules, (3) RadarPLM (LN): fine-tuning only the LN parameters, and (4) RadarPLM (F): full fine-tuning of all model parameters. As shown in the results, the joint fine-tuning of LoRA and LN achieves notably higher performance, while maintaining a relatively shorter training cost (0.116 h/100 epochs) than full fine-tuning (0.148 h/100 epochs). This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
TABLE VIII: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
RadarPLM $\checkmark$ $\checkmark$ 0.116
RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
RadarPLM (LoRA) $\checkmark$ 0.116
RadarPLM (LN) $\checkmark$ 0.112
IV-D3 Loss Function
As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For example, RadarPLM (PA) achieves an average gain of 9.9% over RadarPLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart. In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
TABLE IX: Results of ablation study on the loss function.
@@ IV-D3 Loss Function
TABLE X: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers. The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions).
RadarPLM Performance under Different Configurations
Patch Size Configuration Average DR (%) Average IT (s) Total / Training NP (M)
32 79.44 5.9259 69.39 / 2.46
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
We first investigate the impact of patch size on RadarPLM’s performance. The results of the ablation study are summarized in Table X. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarPLM.
We then investigated the impact of the number of GPT2 layers on RadarPLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table X. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the parameter transfer capabilities of the PLM, while excess layers suffer from overfitting due to the task-irrelevant parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
@@ V Conclusion
In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection. We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization. Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios. These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing. Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.

### v4

@@ III-E Autoencoder-Based Binary Classification Head Retraining
As shown in Fig. 1 and Section III-A, the PLM-based network produces token-level outputs of shape $[B,K,2]$, which is inconvenient for final decision. To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of feature representation $F^{\mathrm{LN}}$. It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector. The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@
3:PLM parameters $\theta_{\text{PLM}}$
@@
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see Eq.(12).
@@
    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see Eq.(14) and Eq.(15).
@@
9:Fine-tuned RadarPLM model with controllable FAR.
@@ IV Experiments
Epochs for RadarPLM fine-tuning 500
@@ IV-A4 Implementation Details
The experimentally relevant hyperparameters and details are shown in Table II. For patching, the size of feature patches is set to 48. For the PLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed. The value of $\alpha$ in Eq.(14) is set to 0.9. All signal feature extraction operations are performed with MATLAB R2016a software. All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
IV-B Performance Comparison
@@
Figure 6: Comparison of model detection performance on various datasets. The accompanying ROC curves demonstrate that the RadarPLM framework achieves the best overall detection performance.
TABLE III: Comparison of detection rates (%) on sufficient training samples under different signal-to-clutter ratio (SCR) conditions.
Method All Data Low SCR Data
$H$ HV VH $V$ AVG $H$ HV VH $V$ AVG
RNN [10] 46.04 46.66 38.18 17.04 36.98 9.45 13.00 11.67 4.47 9.65
Bi-LSTM [10] 70.39 73.15 76.48 60.89 72.48 39.38 42.47 37.48 25.27 36.15
GRU [35] 76.46 77.55 73.96 67.48 73.86 48.99 50.46 49.79 27.33 44.14
Transformer [30] 52.95 71.73 67.32 55.67 61.92 32.31 41.89 46.43 15.35 34.00
PatchTST [36] 75.23 75.72 77.49 66.59 73.76 49.27 49.71 53.18 31.46 45.91
ResNet18 [9] 66.24 80.96 82.31 73.97 75.87 50.45 56.18 53.73 44.54 51.23
OFA [13] 74.19 80.38 79.71 70.36 76.16 49.28 54.18 54.62 26.96 46.26
ADN18 [37] 79.74 83.50 82.01 78.02 80.82 54.76 60.04 31.32 46.11 47.31
RadarPLM 82.51 84.81 84.61 75.90 81.96 60.03 63.19 62.27 44.82 57.58
Note: All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.
TABLE IV: Network parameters (training parameters / total parameters) and average inference time of different methods.
@@
Table III reports the detection rates achieved by RadarPLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarPLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Fig. 6 presents the complete ROC curves; it is evident that RadarPLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarPLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarPLM still delivers a 1.14% gain in the average detection rate.
More Crucially, RadarPLM demonstrates exceptional robustness in the challenging low SCR conditions. As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead. It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%. This result highlights the critical advantage of our method in learning discriminative representation from noisy sequence features.
@@
Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table IV. Although RadarPLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. In particular, RadarPLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
IV-C Scarce Training Sample Condition
@@
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conducted experiments under scarce training sample conditions to assess the effectiveness of RadarPLM. Specifically, we utilized a data split ratio of 10% (Training samples): 5% (Validation samples): 85% (Test samples), where the number of training samples are severely reduced.
@@
In Fig. 7, we present a comparative analysis of all methods under the scarce training sample condition against the standard setting (Abundant Training Samples) described in Section IV-B. The results clearly demonstrate that RadarPLM exhibits the second-lowest performance degradation among all assessed methods, highlighting its robustness to scarce data availability. Specifically, RadarPLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [10]. This marked improvement can be mainly attributed to three key factors:
@@
Figure 7: Performance comparison under scarce and abundant training sample conditions. The results show that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
TABLE V: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
RadarPLM $\checkmark$ $\checkmark$ 0.116
RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
RadarPLM (LoRA) $\checkmark$ 0.116
RadarPLM (LN) $\checkmark$ 0.112
TABLE VI: Results of ablation study on the backbone choice and fine-tuning strategies.
@@
Note: RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results. The best and second-best results are highlighted in red and blue, respectively.
(1). Data1
(2). Data2
(3). Data3
Figure 8: We compare the detection rates achieved by the proposed preference-aware loss against the conventional cross-entropy (CE) loss across various datasets and training epochs. The results demonstrate that while the CE loss exhibits detection rate degradation shortly after a few training epochs (indicating overfitting), the preference-aware loss effectively mitigates overfitting, enabling consistent performance improvement throughout more training epochs.
TABLE VII: Results of ablation study on the loss function.
@@
TABLE VIII: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers.
Parameter Configuration Average DR (%) Average IT (s) Total / Training NP (M)
Patch Size 32 79.44 5.9259 69.39 / 2.46
@@
Note: The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions). The best and second-best results are highlighted in red and blue, respectively.
IV-D Ablation Study
IV-D1 Backbone Architectures
  1. 1.
  2. 2.
  3. 3.
In Table VI, we show the results of the ablation study on backbone choice. It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
IV-D2 Fine-tuning Strategies
To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table V, and the corresponding results are presented in Table VI. In this experiment, four fine-tuning configurations were evaluated: (1) RadarPLM: jointly fine-tuning the LoRA modules and Layer Normalization (LN) parameters, (2) RadarPLM (LoRA): fine-tuning only the LoRA modules, (3) RadarPLM (LN): fine-tuning only the LN parameters, and (4) RadarPLM (F): full fine-tuning of all model parameters. As shown in the results, the joint fine-tuning of LoRA and LN achieves notably higher performance, while maintaining a relatively shorter training cost (0.116 h/100 epochs) than full fine-tuning (0.148 h/100 epochs). This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
IV-D3 Loss Function
As summarized in Table VII, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For example, RadarPLM (PA) achieves an average gain of 9.9% over RadarPLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart. In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
@@ IV-E Analysis of the Effect of Hyperparameters on Experimental Results
We first investigate the impact of patch size on RadarPLM’s performance. The results of the ablation study are summarized in Table VIII. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarPLM.
We then investigated the impact of the number of GPT2 layers on RadarPLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table VIII. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the parameter transfer capabilities of the PLM, while excess layers suffer from overfitting due to the task-irrelevant parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
@@ V Conclusion
In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection. We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization. Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines on low SCR datasets and scarce training data settings. These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing. Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.